OpenStudy (anonymous):

∫[(y^3+5y^2-4)/(y^2),y]

OpenStudy (anonymous):

evaluate the function

OpenStudy (anonymous):

your question is\[\int\limits_{}^{}\frac{y ^{3}+5y ^{2}-4}{y ^{2}}dy\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how to integrate it? please do help me thank you...

OpenStudy (campbell_st):

simplify it 1st \[\int\limits y^3/y^2 dy+ \int\limits 5y^2/y^2dy - \int\limits 4/y^2 dy = \int\limits y dy +\int\limits 5 dy - \int\limits 4y^{-2} dy\] this will give \[y^2/2 + 5y +4y^{-1} + C = y^2/2 + 5y + 4/y + C\]

OpenStudy (anonymous):

oh my God! you're so genius... thank you a lot, OVERWHELMED :)

OpenStudy (anonymous):

i got a correct answer

OpenStudy (campbell_st):

thats ok... with a single term denominator just break it up into smaller problems...

OpenStudy (anonymous):

your'e a blessing to me